Shift summary for Eve, 1/16/00


Subject: Shift summary for Eve, 1/16/00
From: Vivian O'Dell (odell@fnal.gov)
Date: Fri Feb 04 2000 - 15:27:42 CST


 
              Shift Summary for Eve Shift, 1/16/00

On Shift: R. Pordes, A. Tripathi

Hours of beam available at all intensities: 8.5
Hours of beam available at nominal intensity: 8.5
Hours of E799 data (special runs not included): 8.2
Hours of special runs (laser scans, muon runs...):0

Accounting of lost time: A couple of KFRENDS, DDD veto enabled.

========================================================================
Accesses during shift: None

========================================================================
Runs ended this shift:

15544 BEAM_799 19.6 M triggers to tape.

========================================================================
Overview of shift:
  Very smooth data taking.

========================================================================
Description of Shift:

        A couple of kfrends to clear some DYC busies.
        Mounted onsplit tapes UPE022,UPF079,UPK042,UPH059
        Dave N. put 2 new bottles of Ar-Ethane on-line on the west side.

23:50 DDD veto enabled on all planes. Two kfrends.

23:58 DDD veto enabled on ktev1. Ktev1 seems to be processing a lot
        fewer events than other planes. Tapes are almost full anyway, so
        the next shift crew stops the run to do a fresh start.
 
========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift



This archive was generated by hypermail 2b27 : Fri Feb 04 2000 - 15:27:46 CST